# arXiv:2603.00422v1[stat.AP] 28 Feb 2026

### 28 Feb 2026 v1

@@ 4.4 Synthesis
Behavioral, informational, and intervention coupling are distinct mechanisms, but they compound. A platform introduces a policy change (behavioral coupling) that alters supply composition. The demand model does not observe the new supply ceiling (informational coupling). The platform detects the resulting imbalance and intervenes again (intervention coupling), further altering the relationship. The feedback is recursive.
@@
Table 1: Coupling framework: layers, diagnostics, failure modes, and research directions.
@@
Behavioral (Sec. 4.1) Host exit/entry rates; pricing rule changes; cancellation policy shifts Parameters trained pre-policy are invalid post-policy; cross-side propagation missed Elasticity asymmetry diagnostic (§5)
Informational (Sec. 4.2) Search-to-booking gap; occupancy rates; calendar availability Systematic over-prediction when supply binds; censoring unrepresented Moving ceiling; search-to-booking gap; variance decomposition (§5)
Intervention (Sec. 4.3) Regulation dates; platform A/B tests; commission changes Regime-specific parameters applied across regimes; Lucas critique Causal forecasting under unobservable intervention (§5)
@@ 4.5 A simulation-based demonstration
The following toy model illustrates the core failure mode. A market is observed over $T=200$ periods with endogenous supply that responds to lagged demand. Realized bookings are $B_{t}=\min(D_{t},S_{t})$. In the first 150 periods, supply is sufficiently high that the constraint rarely binds. At $t=151$, a regulation shock permanently reduces the supply intercept, simulating a registration requirement that removes casual hosts. Full model specification is given in the Supplementary Material.
@@ 5 Research agenda
The coupling problem implies that forecasting in platform accommodation markets requires methods that jointly model supply and demand, respect endogenous constraints, and remain valid under intervention. While individual building blocks exist in adjacent literatures (revenue management, structural IO, marketplace design), they have not been assembled into an integrated forecasting framework for platform accommodation. Before presenting the open problems, it is useful to map the relevant methodological traditions and their connections to the coupling layers identified above. A more detailed technical survey is provided in the Supplementary Material.
@@ 5.1 Methodological landscape
Four bodies of work bear directly on the coupling problem, each addressing a different facet.
Tourism demand forecasting provides the baseline toolkit: econometric models, exponential smoothing, machine learning, and forecast combination (Song et al., 2019; Song and Li, 2008). This literature excels at modeling demand as a standalone time series and has converged on rigorous evaluation protocols. Its limitation, from the coupling perspective, is that supply enters either as a fixed background condition or not at all. When the supply ceiling is approximately constant, this omission is benign; when it shifts, the toolkit produces forecasts that are predictively adequate in-sample but structurally fragile out-of-sample (Section 4.5).
@@
Causal and intervention inference supplies the tools for evaluating what happens when coupling relationships change. Bayesian structural time series (Brodersen et al., 2015), synthetic control (Abadie et al., 2010, 2015), and difference-in-differences designs are standard for retrospective policy evaluation. The deeper connection is to the Lucas critique (Lucas, 1976): parameters estimated under one policy regime are not invariant to policy change, which is precisely what intervention coupling creates. Extending these tools from retrospective evaluation to prospective forecasting under unobserved intervention is the hardest open problem identified below.
@@
Each tradition solves part of the problem. Tourism forecasting models demand well but ignores the supply constraint. Revenue management handles the constraint but assumes it is fixed. Two-sided economics explains the coupling but does not forecast. Causal inference evaluates past interventions but does not project forward. Compositional methods represent distributional targets but have not been integrated with supply-demand coupling. The remainder of this section identifies five open problems whose resolution would connect these building blocks into an integrated framework. The problems are ordered from the most immediately actionable to the most methodologically challenging.
@@ 5.2 The search-to-booking gap as a supply constraint signal
In accommodation platforms, guests search before they book, and the temporal profile of search activity differs from that of booking activity. The search-to-booking gap is defined here as the divergence between demand intent proxies (searches, listing views, wishlists) and realized bookings, measured at the distributional level across lead times, locations, or price segments. This gap is informative about the difference between latent demand $D_{t}$ and realized bookings $B_{t}$: when the gap widens, it signals that supply constraints, matching frictions, or price misalignment are preventing demand intent from converting to transactions. This signal has not been formalized as a forecasting input. Search volume data are used as leading indicators of aggregate arrivals (Sun et al., 2019; Li et al., 2021), but the informational content of the conversion margin is discarded. The gap can be quantified using compositional distance measures on paired lead-time distributions; see the Supplementary Material for technical details.
@@ 5.3 Demand forecasting under an endogenous supply ceiling
Section 4.2 established that realized demand is bounded by available supply and that this ceiling is time-varying and endogenous. The open problem is to build a demand forecasting framework that explicitly incorporates this constraint.
This has precedent. In labor economics, Barnichon and Nekarda (2012) forecast the unemployment rate by modeling the flows that govern the stock. The simultaneous equation tradition (Haavelmo, 1943; Zellner and Theil, 1962) provides identification strategies for jointly determined systems. A promising direction is a two-equation system where the demand equation conditions on a latent supply state and the supply equation conditions on a demand signal, with the constraint that realized bookings are bounded by both.
The central identification challenge is that $D_{t}$ and $S_{t}$ are not separately identified from $B_{t}\leq\min(D_{t},S_{t})$ without instruments or structural assumptions. Candidate instrument classes include: (i) regulation shocks that constrain supply; (ii) host-side fee or commission changes; (iii) weather or events in origin markets that shift demand; and (iv) timing assumptions exploiting the fact that supply decisions precede demand realization for a given stay date. Each class requires exclusion restrictions that are plausibly violated on platforms. The honest assessment is that clean instruments are rare in this setting, and research designs should state exclusion restrictions explicitly, propose falsification tests, and consider partial identification bounds (Athey and Imbens, 2017).
@@ 5.4 Decomposing demand variance into exogenous and endogenous components
In a coupled system with fast feedback, observed demand variance has two components: genuine exogenous uncertainty and endogenous oscillation from coupling. A testable formulation fits a demand-only model and measures residual variance, then fits a coupled model with supply state variables and measures the reduction. If the reduction is economically meaningful, coupling is a quantitatively important source of recoverable forecast error. The connection to existing methodology is through time-varying volatility models (Katz and Weiss, 2025). The open question is whether volatility clusters are partly endogenous to supply-demand feedback.
@@ 5.5 Price elasticity asymmetry as a coupling diagnostic
Not all markets are equally coupled. When supply and demand respond to price signals with comparable elasticity, the system equilibrates quickly and independent forecasts may suffice. When one side is substantially more elastic, persistent imbalances arise and coupling effects are strongest. A research design would estimate time-varying supply and demand elasticities for a panel of platform markets, construct an asymmetry index, and test whether periods of high asymmetry correspond to periods of poor relative performance of independent forecasts via Diebold and Mariano (1995) tests.
@@ 5.7 Summary and prioritization
The five problems are not independent. The search-to-booking gap provides a measurable mismatch signal. The moving ceiling formalizes this mismatch as a constraint. The variance decomposition quantifies the coupling mechanism’s contribution to forecast error. The elasticity diagnostic identifies when coupling is strongest. Causal forecasting under unobservable intervention addresses the deepest methodological obstacle.
@@
A researcher entering this agenda would do well to begin with the supply measurement problem that underlies several formulations. Without canonical supply state variables that can be measured consistently across markets, the ceiling cannot be estimated, the variance cannot be decomposed, and the coupling cannot be parameterized. The payoff is substantial. Methods developed for platform accommodation will generalize to other platform markets (ride-sharing, labor platforms, e-commerce marketplaces) where similar coupling dynamics are present but less well documented.
5.8 A minimal empirical design
@@
To make the agenda concrete, this subsection sketches a feasible empirical test of the coupling hypothesis. The design requires one city with a well-documented supply shock, one intent proxy, and a standard forecasting evaluation.
@@
Setting. A city that introduced a short-term rental registration requirement with a known effective date. Candidate cities include San Francisco, Barcelona, and Amsterdam, all of which have documented regulation events in the literature (Bibler et al., 2025; Duso et al., 2024; Overwater and Yorke-Smith, 2022).
@@
Supply construction. $S_{t}$ is constructed from daily calendar scrapes as the count of available listing-nights for a 30-day forward stay-date window, following the transparent scrape-based pipeline of Wang et al. (2024).
@@
Evaluation. Two forecasting models are compared out of sample over the post-regulation period: (i) a demand-only AR model trained on pre-regulation $B_{t}$, and (ii) a coupled model that forecasts $\hat{B}_{t}=\min(\hat{D}_{t},\hat{S}_{t})$, where $\hat{S}_{t}$ is a simple inventory nowcast. Accuracy is compared via Diebold and Mariano (1995) tests on MAE. If the coupled model significantly outperforms after the supply shock but not before, this constitutes evidence that coupling matters when regimes change, consistent with the simulation in Section 4.5.
@@
Data access. The supply-side construction and evaluation can be replicated with publicly available calendar data from Inside Airbnb (Inside Airbnb, 2024). A scalar search-to-booking ratio can be constructed from Google Trends (Google, 2024). The full lead-time compositional gap requires platform-internal search logs. If proprietary platform data are available, the same design can be executed with exact booking counts, lead-time-tagged searches, and richer inventory signals; the public-scrape version establishes a replicable baseline.
@@ 6 Implications for destination management and tourism policy
Destination managers and tourism boards. Organizations that forecast visitor volumes to plan infrastructure, staffing, and marketing should recognize that booking data from platform sources reflect supply conditions as much as demand conditions. A city that observes declining Airbnb bookings after introducing a registration requirement is not necessarily experiencing reduced tourism demand; it may be observing the mechanical effect of a supply contraction on a censored outcome. Conflating the two leads to misallocated marketing budgets and incorrect assessments of destination competitiveness. The search-to-booking gap (Section 5.2) offers a practical diagnostic: if search intent remains stable while bookings decline, the binding constraint is supply, not demand.
Regulatory impact assessment. Cities considering short-term rental regulations should anticipate that supply-side interventions propagate to the demand side through behavioral coupling. A registration rule that removes casual hosts does not simply reduce supply by the number of removed listings; it alters the composition of remaining supply (skewing toward professional operators), shifts the price distribution, and changes the demand-side experience. Pre-regulation booking forecasts are unreliable guides to post-regulation outcomes precisely because the coupling relationship has changed. Regulatory impact assessments that use pre-regulation demand trends to project post-regulation bookings will systematically overestimate lost tourism revenue if they do not account for supply-side adjustment.
Platform operators. Forecasting systems that support earnings guidance, treasury operations, and capacity planning are vulnerable to the same coupling failures described here. When platforms modify pricing algorithms, commission structures, or cancellation policies, demand forecasts trained on pre-change data will mis-predict in the post-change environment. Operationally, this means that forecast models should include supply state variables as inputs and that forecast accuracy should be monitored conditional on recent platform interventions, not just unconditionally.

### 19 Apr 2026 v2

@@ 4.4 Synthesis
Behavioral, informational, and intervention coupling are distinct mechanisms, but they compound. A platform or regulator introduces a rule change (intervention coupling). Hosts and guests adjust participation, pricing, availability, and booking behavior (behavioral coupling). The resulting booking series is then observed under a moving and only partly measured supply ceiling (informational coupling). The feedback is recursive, but the decomposition is useful because each layer implies a different empirical problem and a different forecasting failure mode.
@@
Table 2: Coupling framework: layers, proxies, failure modes, and research directions.
@@
Behavioral (Sec. 4.1) Host entry and exit; calendar activation or blocking; price resetting; booking and cancellation responses Cross-side propagation is missed; model parameters fit one participation regime and are carried into another Elasticity asymmetry diagnostic; state-dependent participation modeling
Informational (Sec. 4.2) Search-to-booking divergence; occupancy rates; calendar availability; blocked-booked sensitivity Bookings are misread as unconstrained demand; stock-out censoring is omitted Moving ceiling; search-to-booking gap; variance decomposition
Intervention (Sec. 4.3) Regulation dates; ranking changes; commission changes; cancellation-policy reforms; A/B tests Regime-specific parameters are applied across institutional regimes; Lucas-critique problem Causal forecasting under unobservable intervention
@@ 4.5 A simulation-based demonstration
The following toy example is intentionally stylized. Its purpose is not to establish universal dominance of a coupled model over all possible demand-side benchmarks. It isolates one failure mode: if a model is trained in a regime where supply rarely binds, a later supply shock causes demand-only forecasts to extrapolate a ceiling that no longer exists. A market is observed over $T=200$ periods with endogenous supply that responds to lagged demand. Realized bookings are $B_{t}=\min(D_{t},S_{t})$. In the first 150 periods, supply is sufficiently high that the constraint rarely binds. At $t=151$, a regulation shock permanently reduces the supply intercept, simulating a registration requirement that removes casual hosts. Full model specification is given in the Supplementary Material.
@@
Because the data-generating process is aligned with the coupled model and treats supply as directly observed, this exercise should be read as a mechanism demonstration rather than as a comprehensive horse race. The more demanding empirical question is whether supply-aware models outperform strong demand-only benchmarks when supply is noisy and only partially observed. Section 5.8 therefore proposes an evaluation design with richer demand-only comparators and explicit sensitivity to supply-measurement error.
@@ 5 Research agenda
The coupling problem implies that forecasting in platform accommodation markets requires methods that jointly model supply and demand, respect endogenous constraints, and remain valid under intervention. While individual building blocks exist in adjacent literatures (revenue management, structural IO, marketplace design), they have not been assembled into an integrated forecasting framework for platform accommodation. The agenda below is organized in three layers. Sections 5.2 and 5.3 address foundational measurement and identification problems. Sections 5.4, 5.5, and 5.8 describe near-term empirical designs that can be implemented with currently available public or proprietary data. Section 5.6 is the frontier problem: forecasting under unobserved intervention. Table 4 summarizes the minimum data, identification route, and validation target for each problem. A more detailed technical survey is provided in the Supplementary Material.
Table 4: Research agenda for coupled forecasting in platform accommodation.
Research problem Type Minimum data Plausible design Validation target
Search-to-booking gap Foundational Bookings plus a public intent proxy for the scalar version; internal search logs for the compositional version Scalar ratio or paired-composition distance used as a nowcasting or regime-state variable Gap widens when stock-outs bind and improves forecast performance
Endogenous supply ceiling Foundational Bookings, supply proxy, and either an instrument or a documented supply shock Constrained state-space or simultaneous-equation design with partial identification if needed Lower post-shock bias and coherent latent-demand estimates
Variance decomposition Applied Bookings, supply states, and strong demand-side covariates Nested forecast comparison between demand-only and coupled systems Economically meaningful reduction in MAE, CRPS, or bias attributable to supply states
Elasticity asymmetry Applied Panel of prices, availability, bookings, and exogenous shifters Time-varying supply and demand elasticities linked to forecast losses High asymmetry predicts when decoupled models fail
Unobservable intervention Frontier Long multivariate panels with suspected latent breaks Break-detection or robust-forecasting methods on the coupling relationship itself Calibration is maintained across latent regime changes
@@ 5.1 Methodological landscape
Five bodies of work bear directly on the coupling problem, each addressing a different facet.
Tourism demand forecasting provides the baseline toolkit: econometric models, exponential smoothing, machine learning, and forecast combination (Song et al., 2019; Song and Li, 2008; Song et al., 2023). This literature excels at modeling demand as a standalone time series and has converged on rigorous evaluation protocols. Its limitation, from the coupling perspective, is that supply enters either as a fixed background condition or not at all. When the supply ceiling is approximately constant, this omission is benign; when it shifts, the toolkit produces forecasts that are predictively adequate in-sample but structurally fragile out-of-sample (Section 4.5).
@@
Causal and intervention inference supplies the tools for evaluating what happens when coupling relationships change. Bayesian structural time series (Brodersen et al., 2015), synthetic control (Abadie et al., 2010, 2015), and difference-in-differences designs are standard for retrospective policy evaluation. The deeper connection is to the Lucas critique (Jr, 1976): parameters estimated under one policy regime are not invariant to policy change, which is precisely what intervention coupling creates. Extending these tools from retrospective evaluation to prospective forecasting under unobserved intervention is the hardest open problem identified below.
@@
Each tradition solves part of the problem. Tourism forecasting models demand well but ignores the supply constraint. Revenue management handles the constraint but assumes it is fixed. Two-sided economics explains the coupling but does not forecast. Causal inference evaluates past interventions but does not project forward. Compositional methods represent distributional targets but have not been integrated with supply-demand coupling. The remainder of this section identifies five open problems whose resolution would connect these building blocks into an integrated framework.
@@ 5.2 The search-to-booking gap as a supply constraint signal
In accommodation platforms, guests search before they book, and the temporal profile of search activity differs from that of booking activity. The search-to-booking gap is defined here as the divergence between demand-intent proxies and realized bookings. Two versions should be distinguished. The scalar version compares an aggregate demand-intent proxy, such as Google Trends or card-based travel search activity, with booked nights. The compositional version compares full lead-time, spatial, or segment distributions of search intent and realized bookings.
This distinction matters empirically. Search volume data are already used as leading indicators of aggregate arrivals (Sun et al., 2019; Li et al., 2021; Hu et al., 2025a), but the informational content of the conversion margin is usually discarded. Recent critical assessment of Google Trends and Baidu Index applications in Tourism Management underscores that query selection, preprocessing, and model specification remain central concerns when search-based proxies are used for tourism demand forecasting (Mikulić and Baumgärtner, 2025). From the perspective of this paper, a widening scalar gap can still be useful as a public-data indicator that demand intent is not converting into bookings.
The stronger compositional object requires more data. It can be quantified using paired lead-time distributions and compositional distance measures, but Google Trends does not recover such lead-time distributions by itself. The compositional version therefore requires platform-internal search logs or equivalently granular data in which searches can be linked to intended stay dates. The public-data and internal-data versions are both useful, but they answer different questions and should not be conflated.
@@ 5.3 Demand forecasting under an endogenous supply ceiling
Section 4.2 established that realized demand is bounded by available supply and that this ceiling is time-varying and endogenous. The open problem is to build a demand forecasting framework that explicitly incorporates this constraint while being honest about measurement.
This has precedent. In labor economics, Barnichon and Nekarda (2012) forecast the unemployment rate by modeling the flows that govern the stock. The simultaneous-equation tradition (Haavelmo, 1943; Zellner and Theil, 1962) provides identification strategies for jointly determined systems. A promising direction is a two-equation or state-space system in which the demand equation conditions on a latent supply state and the supply equation conditions on a demand signal, with the constraint that realized bookings are bounded by both. Under platform-internal data, the supply state may be modeled directly from exact calendar and booking states. Under public data, the same problem is better viewed as a proxy or partial-identification problem because the supply variable is measured with error.
The central identification challenge is that $D_{t}$ and $S_{t}$ are not separately identified from $B_{t}\leq\min(D_{t},S_{t})$ without instruments or structural assumptions. Candidate instrument classes include: (i) regulation shocks that constrain supply; (ii) host-side fee or commission changes; (iii) weather or events in origin markets that shift demand; and (iv) timing assumptions exploiting the fact that supply decisions precede demand realization for a given stay date. Each class requires exclusion restrictions that are plausibly violated on platforms. The honest assessment is that clean instruments are rare in this setting, so research designs should state exclusion restrictions explicitly, propose falsification tests, and consider partial-identification bounds (Athey and Imbens, 2017).
A further complication is supply measurement error. With public calendar scrapes, blocked and booked nights are confounded, minimum-stay rules distort night-level availability, and coverage varies across vendors. Practical designs should therefore either (i) model $S_{t}$ as a latent state observed through noisy proxies, or (ii) report sensitivity across alternative scrape-based construction rules and interval bounds. In this setting, an empirical success is not perfect identification. It is a coupled system that reduces post-shock bias, remains coherent with the booking constraint, and is robust to plausible supply-measurement perturbations.
@@ 5.4 Decomposing demand variance into exogenous and endogenous components
In a coupled system with fast feedback, observed demand variance has two components: genuine exogenous uncertainty and endogenous oscillation from coupling. A useful empirical design is therefore nested forecast comparison. Let $M^{D}$ denote a strong demand-only benchmark that uses the best available public demand-side covariates. Let $M^{DS}$ augment that system with supply states or a censoring layer. The variance attributable to coupling is then the reduction in forecast loss or predictive variance when moving from $M^{D}$ to $M^{DS}$, evaluated over rolling windows and especially around supply shocks.
This idea connects naturally to time-varying volatility models (Katz and Weiss, 2026), but the validation criterion should be forecasting, not just model fit. The relevant question is whether adding supply states produces economically meaningful improvements in MAE, MASE, CRPS, post-shock bias, or calibration. With public data, the exercise should be repeated across alternative supply-construction rules so that the variance reduction is reported as robust, attenuated, or fragile to measurement assumptions. That design would quantify whether coupling is merely conceptually plausible or actually a material source of recoverable forecast error.
@@ 5.5 Price elasticity asymmetry as a coupling diagnostic
Not all markets are equally coupled. When supply and demand respond to price signals with comparable elasticity, the system equilibrates more quickly and independent forecasts may suffice. When one side is substantially more elastic, persistent imbalances arise and coupling effects should be strongest. This suggests a diagnostic rather than a universal claim.
A workable research design would estimate time-varying supply and demand elasticities for a panel of platform markets using prices, availability, bookings, and exogenous shifters on each side. The object of interest is an asymmetry index, not the elasticity estimates in isolation. The empirical test is then whether periods and markets with high asymmetry also exhibit poor relative performance of decoupled forecasts, measured for example by Diebold and Mariano (1995) comparisons against coupled alternatives. Validation comes from forecast loss, not just from elasticity significance. This would turn the coupling claim into a conditional statement: coupling matters most where the market fails to re-equilibrate quickly.
@@ 5.7 Summary and prioritization
The five problems are not independent, but they do have a logical progression. The search-to-booking gap and the moving ceiling are foundational measurement and identification problems. Variance decomposition and elasticity asymmetry are near-term empirical designs that can be attempted with current data if the measurement problem is handled carefully. Causal forecasting under unobservable intervention is the frontier problem because it requires methods that remain calibrated when the coupling relationship itself changes.
@@
A researcher entering this agenda would do well to begin with the supply-measurement problem that underlies several formulations. Without canonical supply state variables that can be measured consistently across markets, the ceiling cannot be estimated, the variance cannot be decomposed, and the coupling cannot be parameterized. The payoff is substantial. Methods developed for platform accommodation will generalize to other platform markets (ride-sharing, labor platforms, e-commerce marketplaces) where similar coupling dynamics are present but less well documented.
5.8 A feasible empirical design and benchmark set
@@
To make the agenda concrete, this subsection sketches a feasible empirical test of the coupling hypothesis. The design requires one city with a well-documented supply shock, one public demand-intent proxy, repeated supply measures, and a benchmark set that is stronger than a single univariate AR model.
@@
Setting. A city that introduced a short-term rental registration requirement with a known effective date. Candidate cities include San Francisco, Berlin, and Bordeaux, each of which has a documented regulation episode in the literature (Bibler et al., 2025; Duso et al., 2024; Robertson et al., 2024). Amsterdam remains useful as a modeling case, but Overwater and Yorke-Smith (2022) is an agent-based simulation rather than a documented quasi-experimental regulation study.
@@
Supply construction. $S_{t}$ is constructed from daily calendar scrapes as the count of available listing-nights for a 30-day forward stay-date window, following the transparent scrape-based pipeline of Wang et al. (2024). Because public scrapes cannot cleanly distinguish blocked from booked nights, the analysis should report sensitivity across alternative supply-construction rules or interval bounds rather than rely on a single point estimate.
Demand-intent signals. Publicly available scalar intent proxies can be drawn from Google Trends (Google, 2024), card-transaction data (Grau-Escolano et al., 2026), and online review data (Hu et al., 2022). These variables can support demand-only nowcasting and provide the public-data version of the search-to-booking gap. The full lead-time compositional gap requires platform-internal search logs and is therefore a separate, stronger design.
Benchmark set. The demand-only benchmark set should include at least: (i) a univariate time-series model such as AR or ETS on $B_{t}$; (ii) a demand-only model with public leading indicators; and (iii) a mixed-frequency demand-only model when high-frequency search or review data are available. The coupled model should then be compared against this set, not only against the weakest benchmark.
@@
Evaluation. The coupled system forecasts $\hat{B}_{t}=\min(\hat{D}_{t},\hat{S}_{t})$, where $\hat{D}_{t}$ is allowed the same demand-side information set as the strongest demand-only benchmark and $\hat{S}_{t}$ is an inventory nowcast. Accuracy should be compared out of sample over the post-regulation period using MAE, MASE, post-shock bias, and where possible CRPS or interval calibration. If the coupled model significantly outperforms strong demand-only models after the supply shock but not before, this constitutes evidence that coupling matters when regimes change.
@@
Data access. The public-scrape version establishes a replicable baseline. If proprietary platform data are available, the same design can be executed with exact booking counts, exact calendar states, lead-time-tagged searches, and richer inventory signals. The public and proprietary implementations should be presented as nested versions of the same empirical design, not as interchangeable data environments.
@@ 6 Implications for destination management and tourism policy
Destination managers and tourism boards. The practical task is to distinguish demand weakness from a supply contraction. After a regulation change or platform shock, managers should track booked nights, available inventory, and at least one external demand-intent proxy together. A decline in bookings accompanied by stable demand intent and falling inventory should be treated as a supply shock, not as evidence that destination demand has collapsed. Marketing budgets, staffing plans, and infrastructure decisions should therefore be conditioned on the joint movement of bookings and inventory rather than on bookings alone.
Regulatory impact assessment. Cities considering short-term rental regulation should not infer welfare or tourism effects from raw booking changes alone. A useful post-policy dashboard would separately monitor permit or registration counts, active inventory, host composition, prices, and booked nights. That decomposition allows analysts to distinguish supply withdrawal, price adjustment, and booking conversion effects. Regulatory impact assessments based only on pre-regulation demand trends risk overstating lost tourism revenue because they confound demand changes with mechanically reduced supply.
Platform operators. Forecasting systems that support earnings guidance, treasury operations, and capacity planning should treat major ranking, commission, pricing-policy, or cancellation-policy changes as regime events. In practice this means tagging model vintages by intervention regime, monitoring forecast accuracy conditional on recent interventions, and retraining or switching models when the platform changes the rules of the market. Supply state variables should enter the forecasting system directly, and forecast monitoring should be stratified by inventory tightness rather than assessed only in aggregate.
Researchers using public data. Public-data implementations remain useful, but they should be framed honestly as proxy-based versions of the coupled problem. Calendar scrapes provide noisy supply proxies, and Google Trends provides only scalar intent signals. A defensible public-data application therefore emphasizes sensitivity analysis, benchmark strength, and clear separation between what is identified directly and what is only inferred indirectly.
7 Boundary conditions and limitations
The argument developed here is strongest at the level where the supply constraint is most local: individual listings, narrow segments of close substitutes, or short-run city forecasting around identifiable supply shocks. After broad aggregation or over longer horizons, substitution across neighborhoods and property types, price adjustment, and host entry or re-entry can partially restore market clearing. The claim is therefore not that every market-level booking series is always ceiling-dominated. It is that the risk becomes substantively important when supply changes faster than aggregation can smooth it.
A second limitation is empirical observability. Public calendar scrapes provide imperfect supply proxies because blocked and booked nights are confounded, minimum-stay rules distort night-level availability, and coverage varies across vendors. Listing views, wishlists, and many dashboard metrics are listing-level objects unless the platform aggregates them internally. Public demand-intent proxies such as Google Trends are useful scalar signals, but they do not identify lead-time compositions. Public-data implementations should therefore be treated as proxy-based or partially identified versions of the coupled problem.
A third limitation is that the simulation in Section 4.5 is a stylized mechanism illustration. It isolates a supply-shock failure mode under observed supply and a deliberately simple benchmark. It does not by itself establish the incremental value of coupled forecasting relative to the strongest possible demand-only models with rich leading indicators.
Finally, the research agenda remains prospective. The paper identifies the data structures, identification challenges, and validation standards needed for coupled forecasting, but it does not yet deliver a fully estimated production-ready system. The contribution is therefore conceptual and programmatic, with Section 5.8 intended as a concrete next step rather than as a completed empirical test.